# An Adversarial Non-Autoregressive Model for Text Generation with Incomplete Information

### v2

@@ Abstract
Non-autoregressive models have been widely studied in the Complete Information Scenario (CIS), in which the input has complete information of corresponding output. However, their explorations in the Incomplete Information Scenario (IIS) are extremely limited. Our analyses reveal that the IIS’s incomplete input information will augment the inherent limitations of existing non-autoregressive models trained under Maximum Likelihood Estimation. In this paper, we propose for the IIS an Adversarial Non-autoregressive Transformer (ANT) which has two features: 1) Position-Aware Self-Modulation to provide more reasonable hidden representations, and 2) Dependency Feed Forward Network to strengthen its capacity in dependency modeling. We compare ANT with other mainstream models in the IIS and demonstrate that ANT can achieve comparable performance with much fewer decoding iterations. Furthermore, we show its great potential in various applications like latent interpolation and semi-supervised learning.
Keywords: Non-Autoregressive, GANs, Incomplete Information, Text Generation, Generative Models
@@ 1 Introduction
Figure 1: Comparisons between the CIS and IIS. (a) Translating a sentence. (b) Generating comments based on an emotion label.
Non-autoregressive (NAR) models have lower decoding latency compared with autoregressive (AR) models hence received growing attention from the research community (Huang et al., 2022a). NAR models are emerging in tasks like machine translation (Gu et al., 2018) and text summarization (Liu et al., 2022). These tasks are based on the scenario where the input contains complete information of target sentences. However, another scenario in which the input does not possess complete target information, is seldomly explored in the study of NAR models. For convenience, we denote these two scenarios as the Complete Information Scenario (CIS) and Incomplete Information Scenario (IIS), respectively.
The IIS covers both unconditional and conditional generation tasks, such as generating sentences based on given attributes (Keskar et al., 2019), writing stories based on limited clues (Fan et al., 2018) and supporting semi-supervised learning (Xu et al., 2017). Models for these tasks need to obtain long texts or even infer new data during training process. The decoding latency will be augmented, thereby significantly increasing the computation cost. Extending NAR models (which have lower decoding latency) to the IIS is therefore a promising line of further developments for these tasks.
To this end, we begin with a thorough analysis of existing NAR models which are trained on Maximum Likelihood Estimation (MLE). Due to the lost of word dependencies, these models tend to mix words in different candidates. This problem, which is known as the multi-modality problem (Gu et al., 2018), will be exacerbated in the IIS. As shown in Figure 1, in the CIS, the number of possible candidates and their diversity are significantly constrained by the input so as to maintain balanced input and output information. In the IIS, however, because of the incomplete input information, addition information is needed to obtain the output. Complementing different information will lead to completely different results, so the candidate number and the output diversity will significantly increase. Models without word dependencies will easily mix words in different candidates and generate ungrammatical sentences.
The limitations of existing NAR models can also be analyzed from a theoretical perspective. More specifically, optimizing NAR models with MLE remains non-negative lower bounds between the learned distributions and the real distributions (Huang et al., 2022a). The increase of candidate number in the IIS will increase the lower bounds, and thus increase the discrepancy between the distributions of real data and generated data. This unavoidable limitation will directly lead the models to generate sentences out of real distributions which are always ungrammatical ones.
The inherent limitations of MLE prompt us to pay attention to another family of generative models: Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). The synthetic distributions of GANs can theoretically converge to the real distributions regardless of model structures, so they are free from the multi-modality problem. More importantly, it can obtain high quality samples in one single forward pass, which exactly meets the needs of NAR models. Instead of adopting unstable REINFORCE (Williams, 1992) or biased continuous relaxations  (Jang et al., 2017) to process the non-differentiable sampling operation in language GANs, we follow the research line of representation modeling methods (Ren and Li, 2022) and propose an Adversarial Non-autoregressive Transformer (ANT) for the IIS. There are two features in ANT: Position-Aware Self-modulation for obtaining more reasonable hidden representations; and Dependency Feed Forward network (Dependency FFN) for helping the model to capture more accurate word dependencies in the unstable training of GANs. The experimental results demonstrate that ANT gets comparable performance as existing AR models in the IIS but achieves lower decoding latency. The contributions of this paper can be summarized as follows:
  • We articulate the significance of the IIS in text generation, and compare the differences between the CIS and IIS. Moreover, we reveal the limitations of existing (MLE-based) NAR models in the IIS from both empirical and theoretical aspects.
  • Based on GANs, we propose an Adversarial Non-autoregressive Transformer (ANT), which supports two features, Position-Aware Self-modulation and Dependency FFN, to further improve its performance by providing more reasonable representations and enhance its capacity in dependency modeling.
@@
  • We compare the performance of ANT with existing models. The experimental results demonstrate that ANT can get comparable performance as other models with much lower decoding latency. We also show the great potential of ANT in various applications like latent interpolation and semi-supervised learning. To the best of our knowledge, it is the first work demonstrating the effectiveness of GANs in building NAR models.
This paper is structured as follows. In Section 2, we analyze the limitations of exiting MLE-based NAR models and give an overview to the development of NAR models and language GANs. In Section 3, we introduce the details about ANT. We introduce our experiments and analyze experiment results in Section 4. Finally, we draw a conclusion in Section 5.
@@ 2 Background
2.1 Limitations of MLE in the IIS
NAR models were first proposed for machine translation (Gu et al., 2018) and later extended to other tasks like text summarization (Liu et al., 2022). Most of them are based on MLE and have low decoding latency, but always sacrifice sample quality. Existing NAR models suffer from the multi-modality problem due to the lack of word dependencies. This problem will be augmented in the IIS, since there are more candidates, and it can also be illustrated from a theoretical perspective.
2.1.1 Theoretical Analysis
Given a MLE-based NAR model $P_{\theta}(Y|X)$Huang et al. (2022a) reveal a non-negative lower bound between the leaned distribution and the real distribution. More specifically, $min_{\theta}KL(P_{data}(Y|X)||P_{\theta}(Y|X))\geq C$, where $C=\sum_{i=1}^{l}H_{data}(y_{i}|X)-H_{data}(Y|X)$. $C$ is also known as conditional total correlation. Based on it, we further analyze the difference of the conditional total correlation when the scenario is changed from CIS to IIS. We analyze this problem by considering the transformations from two different input $\mathbb{A}$ and $\mathbb{B}$ to output $\mathbb{T}$. Due to the complexity of real data, our analysis is based on the simplifications in Assumption 1.
Assumption 1.
Each target sequence $Y\in\mathbb{T}$ is consisted of $l$ elements. There are no identical elements in the candidates. Each input $X_{A}\in\mathbb{A}$ has $n$ target sequences as candidates and each input $X_{B}\in\mathbb{B}$ has $m$ target candidates. Each $Y$ can only by mapped by a specific input. All the input and target sequences follow uniform distributions.
Based on Assumption 1, we obtain Theorem 1 to describe the difference of the conditional total correlation between two different input $\mathbb{A}$ and $\mathbb{B}$.
Theorem 1.
The difference between the conditional total correlation of input $\mathbb{A}$ and $\mathbb{B}$ has the following relation: $C_{A}-C_{B}=(l-1)\cdot log(n/m)$.
The proof of Theorem 1 can be found in Appendix A. Theorem 1 reveals the relations between the difference of the conditional total correlation, sequence length $l$, and candidate numbers $n$ and $m$. Furthermore, we obtain a mark on Theorem 1:
Mark 1.
If $n>m$, then $C_{A}>C_{B}$. Higher $l$ and $n$ lead to higher difference between $C_{A}$ and $C_{B}$.
Comparing to the CIS, input in the IIS always corresponds to more possible candidates, so its conditional total correlation will be higher. The larger gap between the learned distributions and the real distributions will lead MLE-based NAR models to generate more ungrammatical results in the IIS.
2.1.2 Existing NAR models in the IIS
To narrow down the gaps between NAR models and AR models, researchers improve performance from different perspectives including simplifying data distributions (Gu et al., 2018), adopting new training objectives (Du et al., 2021), and designing new modeling methods (Ghazvininejad et al., 2019).
Among them, there are two most popular techniques: 1) simplifying output with knowledge distillation (Kim and Rush, 2016); 2) enhancing input based on conditional masked language models (Ghazvininejad et al., 2019). Knowledge distillation simplifies the original one-to-many mapping relations into one-to-one relations, so to prevent models from mixing words in different candidates. However, the IIS requires the model to obtain a number of diverse results with same conditions, so it is necessary to directly model the original one-to-many relations. Conditional masked language models obtain results following an iteration manner, so it will inevitably have higher latency. Although these two techniques are so popular that most of exiting NAR models are deeply bound to at least one of them, they are not applicable methods if we want to build a fully NAR model in the IIS.
Recently, two NAR models, which do not rely on the two techniques above, are proposed: Diffusion-lm (Li et al., 2022) and Directed Acyclic Transformer (DA-Transformer) (Huang et al., 2022b). Diffusion-lm draws the idea of diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021) into text generation. It starts with Gaussian noise vectors and gradually denoises them into specific word representations. Although it is effective on various tasks, it inherits the high latency feature of diffusion models. During inference, it requires 200-2000 steps to obtain the results, which has much higher decoding latency than AR models (Li et al., 2022).
DA-Transformer (Huang et al., 2022b) remedies the multi-modality problem by assigning various candidates to different paths in a directed acyclic graph. Although it obtains remarkable performance in machine translation, its heoretical convergence has not been proved yet. In our experiments, it fails to converge when migrating to the IIS. The key to its successful training is the peak path distributions led by the limited number of candidates. This kind of path distributions can avoid the model changing the candidates of other paths when updating the sample in one path (Huang et al., 2022b). In the IIS, each input corresponds to a lot of diverse candidates, and each vertex thus has many possible next vertices. The model needs to assign comparable transition probabilities to all these vertices and finally leads to a flat path distribution. In this time, the candidates in other paths will be inevitably changed when updating the sample in one path, and the model will finally fail to learn different candidates. Therefore, although there are various NAR models in the CIS, they are not reasonable choices for building fully NAR models in the IIS.
2.2 Language GANs
In contrast, GANs (Goodfellow et al., 2014) indicate a more promising training method for NAR models under the IIS. Their global optimal is achieved if and only if the learned distribution is exactly the same as the target one (Goodfellow et al., 2014). Their effectiveness is demonstrated in various tasks (Karras et al., 2020; Biau et al., 2021; Allouche et al., 2022; Sauer et al., 2022; Bot et al., 2023), but their explorations in NAR text generation are extremely limited. GANs can converge to the global optimal in both AR and NAR structures and thus do not have the multi-modality problem. More importantly, GANs can obtain high quality samples in one single forward pass, which exactly meets the needs of building fully NAR models.
Most of exiting language GANs are based on AR structures (Yu et al., 2017; Che et al., 2017; Lin et al., 2017; Fedus et al., 2018). To tackle the non-differentiable sampling operation, they adopt REINFORCE (Williams, 1992) or continuous relaxations (Jang et al., 2017) which are of either high variance or biased estimators (Lin et al., 2020). Our work thus follows another research line, namely representation modeling method (Ren and Li, 2022). This method firstly maps words into representations and trains a generator to model these representations. It avoids the non-differentiable sampling operation during training process, so the gradients from discriminators can pass through to generators directly. This method is free from the limitations in REINFORCE or continuous relaxations and is demonstrated to be an effective way to train language GANs (Ren and Li, 2022).
@@ 2.2 Language GANs
NAGAN (Huang et al., 2021), which also adopts GANs to build NAR models, is the most similar work to ours, yet our work is different in three aspects: 1) we give in-depth analyses about the limitations of existing MLE-based NAR models from the perspective of incomplete information; 2) the performance of their model is significantly limited by the biased straight-through estimator (Bengio et al., 2013), whereas our model is free from this problem; 3) our proposed facilities: Position-Aware Self-Modulation and Dependency Feed Forward Network can further boost model performance and it is not explored in previous work.
@@ 3.1 Model Structure
Based on the representation modeling framework (Ren and Li, 2022), we propose an Adversarial Non-autoregressive Transformer (ANT) which generates text in a fully NAR manner. As shown in Figure 2, there are three parts in ANT: aligner, discriminator and generator. The aligner maps words into representations, and the generator tries to recover these representations from input. The discriminator needs to identify whether input representations are from the aligner or the generator. We adopt Transformer (Vaswani et al., 2017) as the backbones of all the three parts to support highly parallel computation. An input is firstly added with a positional encoding and fed into encoder layers. Each encoder layer has a multi-head attention (MHA) module and feed forward network (FFN) module. A layer normalization is added after each module.
@@
The aligner is trained to reconstruct words based on the masked input, which is the same as the training process of BERT (Devlin et al., 2019). Following the previous work which adopts representation modeling methods to train language GANs (Ren and Li, 2022), we use the loss function of variational autoencoder (VAE) (Kingma and Welling, 2014) to train the aligner:
@@
A non-autoregressive generator can not adopt previously generated words as the input, so we use trainable representations as input and incorporate latent variables into the representations. The generator then gives output representations $r_{i}$ in different positions and uses the same linear transformation layer $F_{LT}$ in the aligner to transform these representations back into words. The discriminator adopts the output representations from the aligner and the generator ($\mu_{x_{i}}$ and $r_{i}$) as input, and identifies whether they are synthetic or not. Different from image GANs whose discriminators give a single scaler output for an image, our discriminator gives output for each representation. During training, the aligner will be trained first, and its parameters are fixed during the training of the discriminator and the generator. The representations given by the generator need not be transformed into words in training process, so the gradients from the discriminator can directly pass through to the generator.
Casual masks are adopted in both the discriminator and the generator to break the possible symmetry in the input. We use Wasserstein distance (Arjovsky et al., 2017) as the training objective and adopt Lipschitz penalty (Petzka et al., 2018) to regularize the discriminator. More specifically, the loss functions of the discriminator $L_{D}$ and the generator $L_{G}$ are:
$L_{D}=-\mathbb{E}_{\mu_{x_{i}}\sim{P}_{x}}[D(\mu_{x_{i}})]+\mathbb{E}_{z\sim{P}_{z}}[D(G(z))]$ (2)
$L_{G}=-\mathbb{E}_{z\sim{P}_{z}}[D(G(z))]$ (3)
where $\mu_{x_{i}}$ is obtained by the aligner and $z$ is the latent variable sampled from a pre-defined distribution.
@@
However, there is still a gap between our basic model and existing autoregressive models. To improve its performance, we propose Position-Aware Self-Modulation and Dependency Feed Forward Network (Dependency FFN).
3.2 Position-Aware Self-Modulation
An effective sampling method plays a key role in the success of GANs. Recent studies about Transformer in image GANs (Lee et al., 2021) adopt self-modulation (Chen et al., 2019) to sample data. Self-modulation assigns the same shift and scale factors to the normalized results in different positions, which leads the representations in various positions to be highly similar even with positional encodings (as shown in Figure 3 (a)). However, the output of the generator (i.e., word representations in different positions) are of high diversities. Similar input representations can not provide clear signals to the generator to obtain those diverse output representations. It brings difficulties for the generator to converge and recover the data distribution.
@@ 3.2 Position-Aware Self-Modulation
where $z$ is the latent variable, $\mathbf{h}^{\prime}_{i}$ is the hidden representation in the $i$-th position, $MLP(\cdot)$ is a non-linear transformation whose activation function is GELU (Hendrycks and Gimpel, 2016), $LN(\cdot)$ is the layer normalization, $N$ is the length of the sentence, and $\gamma(\cdot)$ and $\beta(\cdot)$ are linear transformations. In Position-Aware Self-Modulation, representations in different positions are calculated based on unique parameters and have clear differences (as shown in Figure 3 (b)), so as to provide more effective signals to recover original sentences.
Our preliminary experiment shows that representations obtained by Position-Aware Self-Modulation are of high diversities, while equipping it in every layer will slow down the convergence. Thus, we adopt it only in the input and use dropout as an additional sampling method by using it in both training and inference stage (Ren and Li, 2022). It can improve model performance by injecting slight random noise and regularizing the model at the same time.

### v3

# Unlocking the Power of GANs in Non-Autoregressive Text Generation

@@ Abstract
Generative Adversarial Networks (GANs) have been studied in text generation to tackle the exposure bias problem. Despite their remarkable development, they adopt autoregressive structures so suffering from high latency in both training and inference stages. Although GANs have potential to support efficient generation by adopting non-autoregressive (NAR) structures, their explorations in NAR models are extremely limited. In this work, we conduct pioneering study of building language GANs based on NAR structures. We identify two issues that constrain the performance of GAN-based NAR models. Firstly, existing methods of incorporating latent variables provide highly similar representations which cannot describe the diversity of different words in sentences. We tackle this problem by proposing Position-Aware Self-Modulation, providing more diverse and effective representations. Secondly, the attention mechanism in Transformer cannot accurately build word dependencies in the unstable training of GANs, and we adopt Dependency Feed Forward Network to enhance the model capacity in dependency modeling. Armed with these two facilities, we propose a GAN-based NAR model, Adversarial Non-autoregressive Transformer (ANT). The experimental results demonstrate that ANT can achieve comparable performance with mainstream models in a single forward pass and has great potential in various applications like latent interpolation and semi-supervised learning.
@@ 1 Introduction
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) denote a powerful family of generative models. Unlike diffusion models (Ho et al., 2020; Saharia et al., 2022) and autoregressive (AR) models (Zhang et al., 2021; Yu et al., 2022) requiring multiple inference steps, GANs can produce high quality samples in a single forward pass, and thus have much lower latency (Sauer et al., 2023; Kang et al., 2023). GANs were first applied to text generation to tackle the notorious exposure bias problem (Bengio et al., 2015), which arises from the discrepancy between training and inference processes in Maximum Likelihood Estimation (MLE)-based AR models. These language GANs (Yu et al., 2017; de Masson d’Autume et al., 2019; Ren & Li, 2023) keep the AR structures and tackle the exposure bias problem by using previously generated words as input in both training and inference stage. When providing consistent training and test processes, they do not support parallel computation and have high latency. The high efficiency nature of image GANs are completely lost in existing language GANs. In theory, the global optimality of GANs is achieved if and only if the learned distributions are exactly same with the real distributions (Goodfellow et al., 2014). More importantly, their convergence does not rely on specific structures. GANs can theoretically obtain high quality samples in a single forward pass. Even so, their explorations in non-autoregressive (NAR) text generation are extremely limited.
In this paper, we unlock the power of GANs in building NAR text generative models. Instead of using AR structures like existing language GANs (Yu et al., 2017; Lin et al., 2017; Che et al., 2017; de Masson d’Autume et al., 2019; Ren & Li, 2023), our model adopts NAR structures supporting high efficiency parallel computation. When benefiting from the generation efficiency, we observe clear gaps between the performance of existing AR language GANs and our NAR language GANs. The main obstacles come from two problems. Firstly, GANs rely on latent variables (which are always from a pre-defined distribution) to support sampling, while existing methods of incorporating latent variables provide highly similar representations. These representations cannot describe the diversity between words and thus leading to inaccurate sentence generation. Secondly, Transformer (Vaswani et al., 2017), widely used in NAR models (Gu et al., 2018; Ghazvininejad et al., 2019), establishes word dependencies solely through the attention mechanism. However, the dynamic weight assignment process becomes unstable during the fragile training of GANs, causing the loss of word dependencies and ultimately resulting in ungrammatical outputs.
Regarding the first problem, we propose Position-Aware Self-Modulation which can provide diverse hidden representations for the model to obtain various words in sentences. For the second problem, we replace the original Feed-forward Network (FFN) module in Transformer to be our proposed Dependency Feed-forward Network (Dependency FFN). Different with the attention mechanism whose dependency is easily lost with poor weight assignment, Dependency FFN provides more stable methods for dependency modeling. Armed with these two facilities, we propose an Adversarial Non-autoregressive Transformer (ANT). The contributions of this work are summarized as follows:
  • We conduct pioneering work of building language GANs in NAR structures. To support the generation of various words in sentences, we propose Position-Aware Self-Modulation which can obtain diverse representations to describe the diversity of different words and improve generation quality. Furthermore, Dependency FFN is proposed to support more stable dependency modeling. Different with the attention mechanism, which is easily influenced by the unstable training process of GANs, Dependency FFN can help the model build more accurate dependencies and thus obtain more grammatical results.
  • Utilizing these two facilities, we propose a GAN-based NAR text generative model—ANT. Existing language GANs employ AR structures, which leads to high latency due to their reliance on previously generated words. ANT, however, generates all words in parallel and support high efficiency generation. It inherits the advantages of GANs, enabling the generation of high-quality samples in a single forward pass.
@@
  • The experimental results demonstrate that ANT achieves performance comparable to existing models, but with significantly lower latency, in both unconditional and conditional generation tasks. Besides, we also explore the potential of ANT in applications like latent interpolation and semi-supervised learning. To the best of our knowledge, it is the first work demonstrating the effectiveness of GANs in building NAR text generative models.
@@ 2 Background
GANs (Goodfellow et al., 2014) are initially transferred to text generation to tackle the notorious exposure bias problem (Bengio et al., 2015). More specifically, existing text generative models adopt autoregressive structures as backbones and use Maximum Likelihood Estimation (MLE) as training objectives. These methods use ground truth as input during training, but reads previously generated words during inference. When the model makes mistakes in generation, these mistakes will be fed into the model as input and the model will be in the state space it has never met during training (Bengio et al., 2015). These mistakes will thus be quickly amplified, leading to a sharp decrease in the quality of the generated samples.
The training of GANs does not need ground truth as input, so they can use generated tokens in both training and inference stage. It tackles the exposure bias problem by providing a consistent generation manner in training and test procedures. In text generation, the generator often models the output word probabilities and sample specific words from these probabilities. This sampling operation, however, is non-differentiable and stops the gradients from being passed through to the generator.
Early study tackles this problem by using either Reinforce (Yu et al., 2017; Lin et al., 2017; Che et al., 2017; Guo et al., 2018; de Masson d’Autume et al., 2019) or continuous relaxations (Nie et al., 2019). However, Reinforce Williams (1992) is in high variance, while continuous relaxations like Gumbel-softmax Jang et al. (2017) are biased estimators. Models based on these two methods rely on pre-training techniques to obtain acceptable performance (Ren & Li, 2023).
@@
Another method is to transform words into representations, and train the generator to obtain these representations. This method avoid the non-differentiable sampling operation during training, so the gradients can be passed to the generator directly. Different with Reinforce and continuous relaxations, which directly model word probabilities, these methods model word representations so they are denoted as representation modeling methods (Ren & Li, 2023). This method can obtain satisfied performance without using any pre-training techniques. We thus build our model based on this method.
NAGAN (Huang et al., 2021), which also adopts GANs to build NAR models, is the most similar work to ours, yet the performance of their model is significantly limited by the biased straight-through estimator (Bengio et al., 2013), whereas our model is free from this problem. Besides, our proposed facilities: Position-Aware Self-Modulation and Dependency Feed Forward Network can further boost model performance and they are not explored in previous work.
Furthermore, existing MLE-based NAR models (Gu et al., 2018; Ghazvininejad et al., 2019) suffer from the multi-modality problem Gu et al. (2018), which tends to mix words in different candidates and obtain ungrammatical results.  Huang et al. (2022a) reveal that the KL divergence between their learned distributions and the real distributions remains non-negative lower bounds. In theory, the learned distribution cannot be exactly same with the real distributions unless words in sentences are independent to each other (which does not match the real situation). Thus, existing NAR models are mainly developed in several specific tasks, while their explorations in more general tasks (like unconditional generation) are extremely limited. This further underscores the pressing necessity of investigating more promising approaches for building NAR models, thereby enabling their application across diverse domains.
@@ 3.1 Model Structure
In this paper, we propose an Adversarial Non-autoregressive Transformer (ANT) which generates text in a fully NAR manner. ANT is based on the representation modeling framework (Ren & Li, 2023). As shown in Figure 1, there are three parts in ANT: Aligner, Discriminator and Generator. The aligner maps words into representations, and the generator tries to recover these representations. The discriminator needs to identify whether input representations are from the aligner or the generator. We adopt Transformer (Vaswani et al., 2017) as the backbones of all the three parts to support highly parallel computation. An input is firstly added with a positional encoding and fed into encoder layers. Each encoder layer has a multi-head attention (MHA) module and feed forward network (FFN) module. A layer normalization is added after each module.
@@
The aligner is trained to reconstruct words based on the masked input, which is the same as the training process of BERT (Devlin et al., 2019). Following the previous work which adopts representation modeling methods to train language GANs (Ren & Li, 2023), we use the loss function of variational autoencoder (VAE) (Kingma & Welling, 2014) to train the aligner:
@@
A non-autoregressive generator cannot input previously generated words, so trainable representations are adopted as input. The generator then gives output representations $r_{i}$ in different positions and uses the same linear transformation layer $F_{LT}$ in the aligner to transform these representations back into words. The discriminator adopts the output representations from the aligner and the generator ($\mu_{x_{i}}$ and $r_{i}$) as input. Different from image GANs whose discriminators give a single scaler output for an image, our discriminator gives output for each representation. During training, the aligner will be trained first, and its parameters are fixed during the training of the discriminator and the generator. The representations given by the generator need not be transformed into words in training process, so the gradients from the discriminator can directly pass through to the generator.
Causal masks are adopted in both the discriminator and the generator to break the possible symmetry in the input. We use Wasserstein distance (Arjovsky et al., 2017) as the training objective and adopt Lipschitz penalty (Petzka et al., 2018) to regularize the discriminator. However, there is still a gap between our basic model and existing autoregressive models and we further propose Position-Aware Self-Modulation and Dependency Feed Forward Network (Dependency FFN) to improve model performance.
3.2 Position-Aware Self-Modulation
An effective sampling method plays a key role in the success of GANs. Transformer based image GANs (Lee et al., 2021) generate different samples by adopting self-modulation (Chen et al., 2019) to incorporate latent variables. Self-modulation assigns the same shift and scale factors to the normalized results in different positions, which leads the representations in various positions to be highly similar even with positional encodings (as shown in Figure 2 (a)). However, the output of the generator (i.e., word representations in different positions) are of high diversities. Similar input representations cannot describe the diversity among different words and thus leading to inaccurate sentence generation.
@@ 3.2 Position-Aware Self-Modulation
where $z$ is the latent variable, $\mathbf{h}^{\prime}_{i}$ is the hidden representation in the $i$-th position, $MLP(\cdot)$ is a non-linear transformation whose activation function is GELU (Hendrycks & Gimpel, 2016), $LN(\cdot)$ is the layer normalization, $N$ is the length of the sentence, and $\gamma(\cdot)$ and $\beta(\cdot)$ are linear transformations. In Position-Aware Self-Modulation, representations in different positions are calculated based on unique parameters and have clear differences (as shown in Figure 2 (b)), so as to provide more effective signals to obtain target sentences.